### v1

# On-Device Self-Supervised Learning of Low-Latency Monocular Depth from Only Events

Abstract
Event cameras provide low-latency perception for only milliwatts of power. This makes them highly suitable for resource-restricted, agile robots such as small flying drones. Self-supervised learning based on contrast maximization holds great potential for event-based robot vision, as it foregoes the need to high-frequency ground truth and allows for online learning in the robot’s operational environment. However, online, onboard learning raises the major challenge of achieving sufficient computational efficiency for real-time learning, while maintaining competitive visual perception performance. In this work, we improve the time and memory efficiency of the contrast maximization learning pipeline. Benchmarking experiments show that the proposed pipeline achieves competitive results with the state of the art on the task of depth estimation from events. Furthermore, we demonstrate the usability of the learned depth for obstacle avoidance through real-world flight experiments. Finally, we compare the performance of different combinations of pre-training and fine-tuning of the depth estimation networks, showing that on-board domain adaptation is feasible given a few minutes of flight.
1 Introduction
Event cameras capture per-pixel changes in brightness at microsecond resolution, while consuming only milliwatts of power [13]. This enables low-latency perception and decision making in environments even when undergoing large visual motion, on platforms that are constrained in power budget, such as small drones.
To make full use of the temporal information in the event stream, the learning pipeline consisting of network architecture and loss function should also operate at high frequency [24]. Ground truth for variables like optical flow or depth is often only available at lower rates of 10-20 Hz [37, 14, 5]. While some datasets allow upsampling of ground truth to higher rates [37, 3], reaching the temporal resolution of an event camera might be difficult and come at the cost of high data rates. This holds back supervised learning at the short time scales perceivable with event cameras.
Figure 1: Self-supervised learning pipeline to estimate dense depth from event-camera data. We improve the time and memory efficiency of the pipeline for enabling online learning onboard of a small drone. This leads to better depth estimates and more successful obstacle avoidance behavior.
Contrast maximization [11] allows self-supervised learning of optical flow, depth and ego-motion from events alone [38, 24, 27, 12]. Because no ground truth is needed, learning and prediction can be done at frequencies of 100 Hz [24] or even 200 Hz [25], with the only limit being the network’s ability to integrate temporal information long enough to capture sufficient motion for learning.
A major advantage of self-supervised learning is that it foregoes the time- and pain-staking process of obtaining ground truth. This enables learning to scale up to enormous datasets. An additional, but typically less emphasized advantage is that self-supervised learning can in principle be performed in the operational environment of a robot or other edge device. Such online, self-supervised learning greatly reduces the need for generalization of the learned model, as training happens directly on data sampled from the test distribution [29]. For visual tasks such as monocular depth estimation, this is particularly important, as generalization of this perceptual capability to different environments than the training environment is notoriously difficult [31, 10]. Online, self-supervised learning does bring additional challenges. The first, main challenge is that not only the network but also the learning framework should be computationally efficient enough to run onboard.
In this work, we improve the efficiency of the pipeline such that on-device learning of low-latency monocular depth and ego-motion becomes feasible. We demonstrate continual learning on-board a small flying drone, and show the usability of the resulting depth for obstacle avoidance. Furthermore, we investigate various combinations of pre-training and on-board learning. When trained on event camera datasets, our small recurrent network shows state-of-the-art depth estimation performance among self-supervised approaches. Our work scratches at the unused potential of on-board, online self-supervised learning, promising smaller reality gaps, leading to better performance.
2 Related work
2.1 (Self-/Un)supervised monocular depth estimation
Self-supervised and unsupervised learning of monocular depth has garnered significant attention since the early works that focused on joint depth-pose estimation for static scenes [36, 30, 16]. These foundational studies have spurred further advancements to handle more complex scenarios, such as dynamic scenes, by integrating optical flow estimation [26, 35], leveraging regularization techniques [21], or incorporating motion segmentation [28]. Additionally, several studies have explored learning camera parameters [17, 6], which is particularly relevant for on-device learning scenarios with unknown cameras.
Wang et al.[33] demonstrate that recurrent networks can enhance depth estimation by effectively utilizing information from multiple frames, resulting in more consistent depth scale predictions. Similarly, Bian et al. [2] introduce a loss term to encourage scale consistency, addressing a critical challenge in depth estimation. Achieving depth predictions with a consistent scale significantly enhances the stability of robot control systems relying on these depth estimates. By combining recurrent architectures and scale-consistent training, our approach aligns with these advancements, offering a robust solution for dynamic, on-device learning during real-world operation.
2.2 Contrast maximization to learn from events
The contrast maximization framework enables the extraction of accurate optical flow information by leveraging the temporal misalignment of accumulated events [12, 11]. This optical flow can be estimated either through model-based methods [27] or using neural networks [38, 18, 24]. The choice of the optical flow model itself offers a spectrum of possibilities, ranging from linear models [38, 18], to segmented representations [24], and even parametrized trajectories [19].
In our work, we adopt the approach outlined in [24], as it aligns with our goal of achieving high-frequency estimation using a neural network. This method is particularly well-suited for handling event trajectories that may exhibit nonlinear patterns, which are common in dynamic and complex scenarios. By using a neural network trained to extract optical flow from event data, we aim to ensure both flexibility and accuracy in environments where traditional linear assumptions might fail.
Moreover, this approach supports efficient inference, which is critical for real-time applications such as robotic navigation or on-device learning. The ability to accommodate nonlinear event trajectories further broadens its applicability, making it a robust choice for extracting motion information from event-based sensors in challenging conditions.
2.3 Self-supervised event-based depth estimation
Early works focused on jointly estimating depth and pose directly from events, employing either contrast maximization techniques [38] or photometric error methods based on event frames [34]. A notable contribution by Zhu et al. was their ability to estimate metric depth through the incorporation of a stereo loss term, enabling absolute depth recovery. These methods demonstrated the potential of event-based sensing for depth and pose estimation in static scenes.
More recent research has expanded these approaches to address dynamic scenes [15], where traditional static-scene assumptions do not hold, and developed more principled frameworks for model-based contrast maximization to jointly estimate depth, ego-motion, and optical flow [27]. These advancements represent a significant step toward understanding and processing complex, real-world motion using event data.
Additionally, some works have explored the integration of intensity images as either inputs or components of the loss function [39]. This hybrid approach leverages the complementary information provided by intensity images to enhance the performance of event-based depth estimation, especially in scenarios where pure event data might lack sufficient structure or texture information.
2.4 On-device (learning of) depth estimation
Already in 2016, Lamers et al. [20] demonstrated the feasibility of learning depth estimation onboard a small flapping-wing drone. Although their approach did not produce dense depth maps, it proved effective for navigation, marking an early milestone in on-device learning for aerial robotics.
More recent advancements have focused on generating dense depth maps onboard drones. Liu et al.[22] developed a system to estimate depth from images for obstacle avoidance on a tiny quadrotor, leveraging recorded real-world datasets for training. Bhattacharya et al. [1], on the other hand, used a simulation-based approach to train depth estimation from events. They successfully transitioned to real-world applications by performing offline fine-tuning on real-world data, enabling effective obstacle avoidance in practice.
In this work, we take a distinct approach by performing fine-tuning onboard the drone in an online fashion during flight, adapting the model in real time while actively avoiding obstacles. This method combines the strengths of self-supervised learning with real-time adaptability, paving the way for more robust and efficient systems capable of handling dynamic environments. By eliminating the reliance on extensive offline fine-tuning or pre-collected datasets, our approach addresses the challenges of real-world deployment more directly.
3 Method
3.1 Optical flow from contrast maximization
Contrast maximization [11, 12] assumes events =$\mathcal{E}=\{e_{k}\}$ are triggered by motion, meaning that a warp $e_{k}=(t_{k},\bm{x}_{k},p_{k})\mapsto e^{\prime}_{k}=(t_{\text{ref}},\bm{x}^{\prime}_{k},p_{k})$ with the correct motion estimate will align it with other events triggered by the same portion of a moving edge, increasing the contrast of the image of warped events (IWE).
We follow the contrast maximization framework as in [24], which estimates optical flow for thin slices of the event stream using a recurrent architecture. By concatenating flows $\bm{u}_{i}$ as $\Delta\bm{x}_{k}=\sum_{i}(\Delta t_{i}\bm{u}_{i})(e_{k})$, events can be warped iteratively to neighbouring slices, with the correct flows leading to sharp IWEs at all reference times along the trajectory:
$\mathcal{L}_{\text{CM}}=\frac{1}{T+1}\sum_{t_{\text{ref}}=0}^{T}\frac{\sum_{k}\bar{t}_{k}(t_{\text{ref}})\kappa(\bm{x}_{k})}{\sum_{k}\kappa(\bm{x}_{k})}$ (1)
where $\bar{t}_{k}(t_{\text{ref}})$ is the timestamp contribution of event $e_{k}$ to the IWE at reference time $t_{\text{ref}}$, and $\kappa$ is a bilinear splatting kernel. We regularize (prevent event collapse) by scaling IWEs by the number of pixels with at least one event and by masking events that get warped out of the image space at any point [18, 24].
3.2 Combining depth and ego-motion into flow
Assuming a static scene and no occlusion/disocclusion, depth and ego-motion can be accurately estimated from monocular video alone [36]. The optical flow used to warp a pixel $\bm{x}$ between different views is constructed from depth $D$ and a camera transformation or relative pose $P$:
$\bm{x}^{\prime}\sim KPD(\bm{x})K^{-1}\bm{x}$ (2)
with $K$ the camera intrinsic matrix, $P$ consisting of a rotation $R$ and a translation $\bm{t}$, and $\sim$ because depth is only defined up to a scale. [^1] For simplicity, we omit the conversion to homogeneous coordinates. The network estimates depth $D$ directly using a softplus activation [17]; relative pose $P$ is estimated with rotation expressed in exponential coordinates $\bm{\omega}$ [32], and converted to $R$ using Rodrigues’ formula.
To encourage consistent scale for consecutive depth predictions, we include the geometry consistency loss from [2], which computes a normalized difference between the forward-projected depth $D_{0\shortrightarrow 1}$ and interpolated depth $D^{\prime}_{1}$ for all valid pixels $\bm{x}\in V$ (visible in both images):
$\mathcal{L}_{\text{geo}}=\frac{1}{\lvert V\rvert}\sum_{\bm{x}\in V}\frac{\lvert D_{0\shortrightarrow 1}(\bm{x})-D^{\prime}_{1}(\bm{x})\rvert}{D_{0\shortrightarrow 1}(\bm{x})+D^{\prime}_{1}(\bm{x})}$ (3)
where we average over the number of valid pixels $\lvert V\rvert$. Setting an appropriate weight $\lambda$, this then results in the full loss formulation as:
$\mathcal{L}=\mathcal{L}_{\text{CM}}+\lambda\mathcal{L}_{\text{geo}}$ (4)
3.3 Optimizations for on-device learning
For efficient prediction, we make use of a relatively small, 430k-parameter network. It consists of a strided convolutional encoder, a ConvGRU recurrent bottleneck, and a two-tailed convolutional decoder for depth and ego-motion (more details in the supplementary material). While one network forward pass takes less than a millisecond on an NVIDIA RTX 4090, computing the contrast maximization loss and backpropagating the resulting gradients each take up more than 10 ms.
To make on-device learning feasible, we have to improve the efficiency of the components that make up the loss computation and network update: i) warping all events in the accumulated set of events $\mathcal{E}$ using a sampled optical flow to all reference times $t_{\text{ref}}\in[0,T]$, ii) splatting them to the IWE at that $t_{\text{ref}}$, iii) computing the gradient with respect to the network parameters.
Previous work [24] warped and splatted events in batches using PyTorch functions. This has multiple inefficiencies. Batching different amounts of events together leads to padding with zeros, resulting in wasteful computation and memory usage. This also goes for warping events that already went out of the image space (and therefore do not contribute to the loss anymore). Furthermore, PyTorch does not have an optimized implementation for bilinear splatting, meaning you have to implement it yourself. This results in extra computations and memory consumption for things like automatic differentiation.
Figure 2: Runtime and peak increase in memory consumption for different phases of computing the contrast maximization loss on an NVIDIA RTX 4090 and Jetson Orin NX (log-scale). Naive PyTorch processes all events in a for-loop. While batching events together improves a lot over this, parallel processing of all events in CUDA results in even bigger speedups with less memory consumed.
The abovementioned issues can be resolved by considering that all events independently contribute to the loss because they are summed in the IWEs, and that we can therefore parallelize over all the accumulated events $\mathcal{E}$. We implement the functions to do so in CUDA, and connect them to PyTorch as an extension. As is shown in Fig. 2, the resulting improvements are, depending on the device, roughly 100x in terms of runtime. Common datasets like UZH-FPV [9] and MVSEC [37] have between 1k and 10k events per bin, well within the range of these speedups. Furthermore, looking at the peak delta memory, we can see that not needing padding allows much lower memory consumption when event data is highly sparse. Because we now provide the analytical gradients in the CUDA backward kernel, these do not have to be computed through automatic differentiation, leading to efficiency improvements for the backward pass.
3.4 Using depth for obstacle avoidance
In the absence of metric depth and with a possibly varying scale, we can construct simple obstacle-avoiding behaviour using the difference in predicted depths for different parts of the field-of-view [22, 4]. More specifically, we slice the depth map into $K$ vertical bins, compute the average inverse depth $d_{k}$ for each, and use these to set a desired yaw rate $\dot{\psi}$:
$\dot{\psi}$ $=\dot{\psi}_{\text{goal}}(\bm{d})+\dot{\psi}_{\text{avoid}}(\bm{d})$ (5)
$\dot{\psi}_{\text{goal}}(\bm{d})$ $=\lambda_{\text{goal}}(\operatorname*{arg\,min}_{k}(d_{k})-\bar{k})$ (6)
$\dot{\psi}_{\text{avoid}}(\bm{d})$ $=\lambda_{\text{avoid}}\sum_{k=0}^{K-1}(\bar{k}-k)e^{(-\frac{\alpha}{d_{k}})}e^{(-\frac{(k-\bar{k})^{2}}{2\sigma^{2}})}$ (7)
where yawing to the right is positive, and $\bar{k}=\frac{K-1}{2}$ is the center index. The resulting behaviour is both obstacle-avoiding (avoid part) and depth-seeking (goal part).
Figure 3: Overview of the drone (left) and the flight environment (right). System components in blue are for the on-board depth learning pipeline, orange components are for low-level flight control, and green components are for logging only.
3.5 Robot
The experimental setup as shown in Fig. 3 consists of a custom 5-inch quadrotor with a total weight of approximately 800 g, including all sensors, actuators, on-board compute, and battery. All algorithms are implemented to run entirely onboard, leveraging an NVIDIA Jetson Orin NX embedded GPU to obtain data from the event-camera, perform learning and estimate depth in real-time. Control commands (yaw rate) based on the predicted depth maps are sent to the flight controller, a Kakute H7 mini running the open-source autopilot software PX4.222https://px4.io/ Communication between PX4 and the Orin is done using ROS2 [23]. An MTF-01 optic flow sensor and rangefinder enable stable autonomous flight where only yaw rate is controlled by the depth pipeline.
To keep the event rate down (around 1 Mev/s), we only turn on every fourth pixel on the DVXplorer Micro, resulting in a 160x120 stream (instead of 640x480) for the same field-of-view. These events are accumulated into 20 ms windows and made into a frame for the network. The whole events-to-depth pipeline is running at approximately 30 Hz while learning, consuming on average around 15 W. We include a RealSense D435i depth camera for logging purposes only. To plot ground-truth flight trajectories, we record the drone’s position using a motion capture system.
4 Experiments
4.1 Benchmarks
Figure 4: Qualitative results of disparity predictions on the DSEC disparity benchmark. Images are for visualization only, as disparity estimation is event-based. The same color map is applied to the disparity values from the stereo- and supervised-learning-based method from Cho et al. [8] and ours for easy comparison.
We train our proposed network on the training sets (detailed training sequences provided in the supplementary materials) with a batch size of 8 and a constant learning rate of 1e-4 with the Adam optimizer for 50 epochs. The quantitative evaluations on MVSEC and DSEC are provided in Tab. 1 and Tab. 2.
outdoor_day1 outdoor_night1
Depth Cutoff 10m 20m 30m 10m 20m 30m
Zhu et al. [39]  [^3] The network uses events as input but was trained with intensity frames in the loss function. Therefore, it is included as a reference but is not directly comparable to self-supervised methods trained solely on event data. 1.40 2.07 2.65 2.18 2.70 3.64
Zhu et al. [39] 3.90 3.79 4.89 5.55 4.57 5.72
Zhu et al. [38] 2.72 3.84 4.40 3.13 4.02 4.89
Ours 2.25 3.36 4.23 3.25 3.83 4.50
[…]
Table 1: MAE (mean absolute error) of depth prediction in meters on MVSEC test sequences at various depth cutoff distances. The best result is highlighted in bold, and the second best is underlined. The method shown in the shaded row serves as a reference and is not directly comparable to the others, as it also uses image frames.
1PE 2PE MAE RMSE
SL Cho et al. [8] 8.966 2.345 0.501 1.175
DSEC Baseline [14] 10.92 2.905 0.576 1.381
[…]
Ours (approx. scale) 84.92 70.47 4.946 6.274
Table 2: Quantitative evaluation on DSEC disparity benchmark. Due to the lack of other SSL (self-supervised learning) methods on the leaderboard, we compare against two representative SL (supervised learning) methods.
On MVSEC, our method outperforms the other two self-supervised, events-only baselines [38, 39]. For context, we also provide the results from the approach in [39], which additionally uses intensity frames in the training process for a photometric consistency loss. Although it achieves a higher accuracy, our networks rely solely on event streams during training. For completeness, we also assess the accuracy of dense depth (i.e., not masked by events), as shown in the last row of Tab. 1.
In the absence of self-supervised methods on the DSEC disparity benchmark, we compare our approach against two top-performing stereo-event-based supervised learning baselines [7, 14]. To convert our monocular unnormalized depth predictions from our network output into metric depth, we apply a scaling factor derived from the ratio of the median predicted depth to the ground truth median from the training set, labeled as “approx. scale” in Tab. 2. Additionally, we conduct a grid search on the scaling factor to achieve the highest accuracy on the test set, reported as “best scale”.
While our accuracy on the DSEC disparity benchmark falls short of supervised baselines, qualitative comparisons in Fig. 4 demonstrate that our approach effectively captures meaningful structures within disparity maps, even without ground truth labels during training. Notably, close objects—such as the car in interlaken_00_a(540) and traffic signs in thun_01_b(400), interlaken_01_a(1680), zurich_city_12_a(400) are accurately represented. Although the boundaries in our results may lack the sharpness achieved by supervised baselines, our approach better preserves contour shapes, such as the front of the car in interlaken_00_a(540), the arc of the tunnel in interlaken_00_b(560) and the pole in zurich_city_13_a(260). This advantage is especially evident for thin objects like the sign pole in interlaken_01_a(1680) and zurich_city_12_a(440), which are often challenging for supervised methods to capture accurately. Additionally, our self-supervised approach is immune to artifacts typically caused by the sporadic availability of ground truth at the image frame boundaries, resulting in smoother disparity maps free from discontinuity artifacts.
Several factors constrain the accuracy of our methods on these benchmarks, including the reliance on self-supervised learning with events only, a compact network architecture, the use of monocular depth estimation rather than stereo and the imperfect estimation of scaling factor for converting monocular depth to metric depth. While our aim is not to surpass state-of-the-art methods, we believe the quality of our depth predictions is sufficient to support downstream tasks like robot navigation.
4.2 Robotic Experiments
We first pretrain our network on the UZH-FPV dataset [9] using our self-supervised pipeline. This dataset was chosen for its diverse set of motion trajectories, enabling the network to learn a latent representation that generalizes well across various motion types. After pretraining, the network is deployed on a drone, where online learning (fine-tuning) is performed during flight. The network’s forward pass operates at an average speed of 30 Hz, with a backward pass and gradient update conducted every 10 forward passes. Following each gradient update, the state of the recurrent network is detached from the computational graph without resetting it. This approach ensures bounded memory usage, mitigates potential issues with gradient explosion or vanishing, and allows the network to retain temporal context effectively.
(a) Initialize from a network with random weights. Intervention rate: 50%.
(b) Initialize from pre-trained network, disable online learning. Intervention rate: 35%.
(c) Initialize from pre-trained network, enable online learning. Intervention rate: 11%.
Figure 5: Top-view flight trajectories from various experiments. Blue trajectories represent autonomous flights, while orange trajectories indicate manual flights where human intervention was necessary to prevent potential collisions. The setup with online learning demonstrates the best obstacle avoidance performance, achieving the lowest intervention rate.
Figure 6: Qualitative visualization of disparity map evolution during online learning. Note that the image is for visualization purposes only, as disparity estimation is event-based. The same colour map is applied to predictions from all different models for easy comparison. The pre-trained network begins at step 0, followed by 12K steps of online learning (OL) with streaming event data during flight. For comparison, we also show the prediction quality of a randomly initialized network trained from scratch (TFS) for 12K steps.
We show the metric improvements achieved during online learning in Fig. 7, where saved checkpoints are evaluated on a test sequence recorded in the same environment but with different placements for obstacles. The model shows significant improvement not only in the RSAT (ratio of squared average timestamps) metric, which is strongly correlated with the contrast maximization loss used to optimize the network, but also in MAE (mean absolute error) when compared against disparity ground truth. Additionally, the fine-tuning process is efficient, converging within just a few minutes of flight.
Qualitative results at different snapshots during online learning are presented in Fig. 6. Compared to step 0 (the pretrained model), the disparity values for certain close objects, such as the wall and poles, increase, as evidenced by the brightened colors in those regions. To highlight the advantages of pretraining, we also compare our model with a network initialized with random weights and trained using the same amount of online learning data, as shown in the last row of Fig. 6. The trained-from-scratch network fails to produce meaningful disparity maps within the flight’s limited timespan, underscoring the importance of pretraining for achieving fast and reliable adaptation during online learning.
Figure 7: Mean absolute error (MAE) of disparity prediction and the ratio of squared average timestamp (RSAT) during online learning in flight. Model checkpoints were saved periodically and evaluated on a test sequence unseen by the model beforehand. 300 learning steps correspond to roughly 100 seconds of training during flight.
Finally, we plot the trajectories of the drone during the flight experiments under different settings in Fig. 5. During each experiment, the drone takes off and immediately starts to fly autonomously. A human pilot monitors the flight and intervenes if the drone is in a near collision with an obstacle. After the human pilot corrects for the collision course, the drone is switched to autonomous flying again. When the drone was initialized with random weights, the drone does not avoid obstacles and mostly fly with straight lines, as shown in Fig. 5(a). This resulted in constant human interventions and a highest intervention rate of 50 %, determined by the ratio of time of human piloting during a flight.
When we start the experiment with a pre-trained network, the obstacle avoidance performance improves dramatically, lowering the intervention rate to 35 % in Fig. 5(b). While the drone was already flying around obstacles, fine-tuning the model during flight further lowers the intervention rate to 11 %, as shown in Fig. 5(c), allowing the drone to smoothly fly between obstacles with more curvy trajectories than in previous other settings.
5 Conclusion
We have improved the efficiency of self-supervised learning of monocular depth estimation from camera events, such that on-device learning of low-latency monocular depth and ego-motion becomes feasible. The proposed approach features more efficient splatting, and has been implemented in CUDA instead of PyTorch. For common event rates (e.g., 10e6 Ev/s), it uses 100$\times$ less memory and computational time. When trained and benchmarked on event camera datasets, our small recurrent network outperforms other self-supervised approaches on the MVSEC dataset and captures essential structures with sufficient quality to support downstream navigation tasks on the DSEC dataset. Furthermore, we demonstrate online learning on-board a small flying drone, and show the usability of the resulting depth for obstacle avoidance. Finally, we investigate various combinations of pre-training and on-board learning, showing that the latter leads to a flight intervention rate reduction from 35% to 11%.
Our work scratches at the surface of the unused potential of on-board, online self-supervised learning. The current results already demonstrate that online learning leads to better performance in the operational environment. This confirms that learning with samples from the test environment will be important for challenging vision tasks such as monocular depth estimation from events. Moreover, it is important to note that the performance gap between self-supervised learning and supervised learning is still substantial (see Tab. 2). This suggests that there is ample room for improvement of the self-supervised learning scheme, which may lead to a larger performance improvement after online refinement. Moreover, improved detail in the estimated depth maps will facilitate more complex obstacle avoidance and navigation behaviors. Future work under such conditions would be even better positioned to investigate the downstream behavioral advantages of online self-supervised depth learning.
Self-supervised learning has turned out to be instrumental for learning advanced AI methods on enormous, unlabelled datasets. The current work is a small step towards unlocking this type of learning’s second major advantage, i.e., strongly reducing the need for generalization by learning in the test environment. We expect this to form a key component of truly autonomous robots that can quickly and aptly adapt to their environment.
Supplementary Material
6 Optimized warping and splatting
Fig. 8 visualizes, on a high level, how our optimizations with respect to warping sets of events to different reference times are reducing runtime and memory consumption by preventing unnecessary warping operations.
(a) Warping events to multiple reference times in a batched manner involves zero-padding all bins to the same length (light-gray extensions), leading to unnecessary warping operations. Furthermore, events that were already warped out of the image space (and can be discarded) will still be warped.
(b) Warping all events individually in parallel does away with padding and allows to warp only those events that are still in the image space, giving the minimal amount of warping operations necessary.
Figure 8: Comparison of warping three bins of events to three reference times. Events originating from different bins have different colors. At every stage of warping (1, 2), events get warped to a neighboring bin, until we have reached all three reference times.
7 Extra DSEC qualitative results
We present additional qualitative results from the DSEC test set in Fig. 9. While the self-supervised results exhibit less sharp boundaries, they are free from many artifacts commonly introduced by supervised learning, such as discontinuities at frame boundaries and difficulties in handling thin objects.
8 More robot experiment results
Figs. 10, 11, 12 and 13 show extended qualitative results for four unseen scenes during a flight experiment. Looking at Fig. 10, we see that network’s ability to maximize the contrast of the image of warped events increases quickly from ‘step 0’ (only pre-training) to ‘step 3k’ (pre-training + 100 seconds of learning). While this improvement in terms of contrast maximization loss may be mostly due to just learning to increase the magnitude of the optical flow (as shown by the change from orange to purple in column 3, and black to red in column 5) through scaling depth and ego-motion, subsequent learning steps improve the depth map in more sophisticated ways, judging from the disappearance of the wrong ‘depth gap’ in the center of the disparity images in the third and fourth column. Similar priorities in learning (first scaling, then the rest) can be seen in the other scenes.
9 Training details
For offline training on datasets, we train for 50 epochs with the Adam optimizer and a learning rate of 1e-4. For contrast maximization, we accumulate 10 bins of events, and warp all events to all bin edges. Furthermore, we set the weight for the geometric consistency loss $\lambda=0.05$. For on-device learning, we lower the learning rate to 1e-5. Other specifics vary per dataset, and are mentioned below. In all cases, event streams are rectified.
For MVSEC, we train on outdoor_day2 with input bins of 20 ms of events. We use a batch size of 8, and augment the data with polarity and left-right flips. Training takes around 50 minutes on an RTX 4090.
[…]
10 Network details
We make use of a small convolutional recurrent network to predict depth and ego-motion. The encoder and memory backbone are shared between the depth and ego-motion decoders. Tab. 3 lists the details per layer. In addition, we make use of ELU activations as we experienced dying ReLUs. Also, to prevent border artifacts, we use reflect padding for all convolutional layers.
Table 3: Network layer details. Unless specified otherwise, we use ELU activations, and a kernel size of 3, biases and ‘reflect’ padding for convolutional layers. Total parameter count is 430,416.
Layer type Input shape Output shape # Parameters
Encoder Conv2D(ksize=7, stride=2) $(2,H,W)$ $(16,H/2,W/2)$ 1,584
ResidualConv2D(stride=2) $(16,H/2,W/2)$ $(32,H/4,W/4)$ 18,528
ResidualConv2D(stride=2) $(32,H/4,W/4)$ $(64,H/8,W/8)$ 73,920
Memory ConvGRU $(64+64,H/8,W/8)$ $(64,H/8,W/8)$ 221,568
Depth Conv2D $(64,H/8,W/8)$ $(64,H/8,W/8)$ 36,928
Conv2D(bias=False) w/ SoftPlus $(64,H/8,W/8)$ $(1,H/8,W/8)$ 576
Upsample(scale=8, ‘bilinear’) $(1,H/8,W/8)$ $(1,H,W)$
Ego-motion Conv2D(stride=2) $(64,H/8,W/8)$ $(64,H/16,W/16)$ 36,928
Conv2D(stride=2) $(64,H/16,W/16)$ $(64,H/32,W/32)$ 36,928
Conv2D(bias=False) w/ Identity $(64,H/32,W/32)$ $(6,H/32,W/32)$ 3,456
AdaptiveAvgPool2D $(6,H/32,W/32)$ $(6,1,1)$
11 Robot details
We built a 5-inch quadrotor for our robotic experiments. The drone is equipped with on-board sensors that provide the flight controller with all relevant information to follow high-level control commands. More specifically, the EKF running on the Kakute H7 Mini flight controller fuses IMU measurements with velocity and height measurements coming from a MTF-01 optical flow/range sensor into a stable position and velocity estimate. This allows a neural network (like our depth network) to give high-level commands like velocity setpoints or rotational rates.
All relevant components on the drone, along with their weight and power consumption, can be found in Tab. 4. All separate modules on the Jetson Orin NX are connected via ROS2, which also allows for logging to rosbags. ROS2 can also be connected via UART to the internal publish-subscribe messaging API that is used in the PX4 flight controller firmware.
The power consumption during flight is measured by keeping track of the total mAh consumed by over two flight tests. Together with the measured power consumption of the Jetson using jtop and the expected maximum power draw of both cameras, this allows us to calculate the power consumption of the drone (see Tab. 4).
Table 4: List of hardware components used during robotic experiments. Power consumption estimates are obtained from Jetson’s jtop, battery drain during flight experiments, or component datasheets.
Component Product Mass [g] $\bm{\sim}$Power [W]
Frame Armattan Marmotte 5 inch 455 200
Motors Emax Eco II Series 2306
Propellers Ethix S5 5 inch
Flight controller Holybro Kakute H7 Mini
Optical flow & range sensor MicoAir MTF-01
ESC Holybro Tekko32 F4 4in1 mini 50A BL32
Battery iFlight Fullsend 4S 3000mAh Li-Ion 208 -
On-board compute NVIDIA Jetson Orin NX 16GB & DAMIAO v1.1 carrier board 62 9
Event camera iniVation DVXplorer Micro 22 max 0.7
Stereo camera Intel RealSense D435i 75 max 3.5
Total - 822 213.2
Figure 9: Additional qualitative results of disparity predictions on the DSEC disparity benchmark. Images are for visualization only, as disparity estimation is event-based. The same color map is applied to the disparity values from the stereo- and supervised-learning-based method from Cho et al. [8] and ours for easy comparison.
Figure 10: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (‘step 0’), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (colormap shared between rows) and local (colormap for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local colormaps, respectively.
Figure 11: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (‘step 0’), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (colormap shared between rows) and local (colormap for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local colormaps, respectively.
Figure 12: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (‘step 0’), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (colormap shared between rows) and local (colormap for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local colormaps, respectively.
Figure 13: Extended qualitative results on unseen data from a flight test recording. From top to bottom, we evaluate a pre-trained-only network (‘step 0’), then four networks after increasing amounts of online learning (OL), and finally a network trained-from-scratch. The bottom row starts with the single 20 ms bin of events currently seen by the network. The rest of the second column shows the accumulated events (multiple bins) in the current contrast maximization loss window. Applying the iterative warp by the optical flow constructed from depth and ego-motion gives the warped and deblurred events in the first column. Columns three and four show disparity at a global (colormap shared between rows) and local (colormap for only that row) scale. The last two columns show optical flow constructed from depth and ego-motion with global and local colormaps, respectively.